On January 1, 2017, Arbor Corp issued $800,000 of 20-year, 11% bonds at market yield of 12% p.a. Interest is payable semiannually on June 30 and December 31. Compute the issue price of the bonds. Show the financial statements effects using the template for the following: 1) bond issuance, 2) semiannual interest payment and discount amortization on June 30, 2017, and 3) semiannual interest payment and discount amortization on December 31, 2017 Price of a bond is $739,814.81 How do you record this on a financial effects table > Cash+NonCash=Liab+Contributed Capital+Earned Capital
Answer #1
Asset = Liabilities + equity
cash + Non cash = Liabilities + contributed capital + Earned capital
1 739814.81 739814.81
2 -44000

388.89

-44388.89
3 -44000 412.22 -44412.22

Working:

Semiannual interest paid =par value *interest rate *n/12

                      = 800000 * .11*6/12

                       = 44000

Date Interest expense Interest paid Discount amortized carrying value at end
1/1/2017 739814.81
30/6/2017 739814.81*.12*6/12 = 44388.89 44000 388.89 739814.81+388.89= 740203.70
31/12/2017 740203.70*.12*6/12= 44412.22 44000 412.22 740203.70+412.22= 740615.92